OpenStudy (anonymous):

how to solve this indice simultaneous eqn? 3^x-1 = 9x3^y ------- (1) 2^x + 2^y = 9 ------- (2)

OpenStudy (lalaly):

is that (9x^3)^y?

OpenStudy (anonymous):

its 9 times 3^y.

OpenStudy (lalaly):

oh okk

OpenStudy (anonymous):

thanks :)

OpenStudy (lalaly):

and is that 3^x -1 or 3^(x-1)

OpenStudy (anonymous):

3 to the power of (x-1).

OpenStudy (anonymous):

change it so it all has the same base 3^(x-1) = 3^3y

OpenStudy (lalaly):

ok then the first equation becomes 3^(x-1) = 3^(2+y)

OpenStudy (anonymous):

ok.i got the first one.but the second equation has different bases of 2 and 9.

OpenStudy (anonymous):

actually i am stuck in the second equation about how to change it to same base.

OpenStudy (anonymous):

\[2^{x} + 2^{y} = 9\]

OpenStudy (lalaly):

x-1=2+y x=3+y nw substitute 2^(2+y) + 2^y =9 4*2^y +2^y =9 2^y(1+4)=9 2^y=9/5

OpenStudy (anonymous):

this is where i got stuck.

OpenStudy (anonymous):

y=0. x=3

OpenStudy (anonymous):

thats the answer written in the book.i just need to know how to solve it.

OpenStudy (anonymous):

right. x=3+y. put this in the second equation. so 2^(3+y)+2^y=9. this is 9.2^y=9. i.e 2^y=1. so y=0. this gives x=3, sonce x=3+y

OpenStudy (anonymous):

*since

OpenStudy (anonymous):

this is 9.2^y=9. i.e 2^y=1. so y=0.. can you explain this part bro?

OpenStudy (anonymous):

i understand the substitution part.but not what happens afterwards.

OpenStudy (anonymous):

2^3=8. so 2^(3+y) + 2^y =9 simplifies to 8.2^y + 2^y =9. this reduces to 9. 2^y =9. get it?

OpenStudy (anonymous):

bro i am really sorry but u mind explaining how you get it again?really appreciate it man.

OpenStudy (anonymous):

2^3 = 8 but how come it becomes 8.2?

OpenStudy (anonymous):

fine, ill walk you through it. 3\[3^{x-1}=3^{2+y}\] right? equate the exponents.. thsi gives \[x=3+y\] put this in the second equation \[2^{3+y} + 2^ y =9\] this is \[8\times 2^y + 2^y = 9\] which is \[9 \times @^y = 9\] y=0. x=3

OpenStudy (anonymous):

OH I GOT IT BRO.I just got confused by where 8.2 came from.

OpenStudy (anonymous):

Really a great help man..Wish you were my math teacher. :)

OpenStudy (anonymous):

Thanks again bro :)

OpenStudy (anonymous):

see ya dude :)